### v1

# LACONIC: A 3D Layout Adapter for Controllable Image Creation

Abstract
Existing generative approaches for guided image synthesis of multi-object scenes typically rely on 2D controls in the image or text space. As a result, these methods struggle to maintain and respect consistent three-dimensional geometric structure, underlying the scene. In this paper, we propose a novel conditioning approach, training method and adapter network that can be plugged into pretrained text-to-image diffusion models. Our approach provides a way to endow such models with 3D-awareness, while leveraging their rich prior knowledge. Our method supports camera control, conditioning on explicit 3D geometries and, for the first time, accounts for the entire context of a scene, i.e., both on and off-screen items, to synthesize plausible and semantically rich images. Despite its multi-modal nature, our model is lightweight, requires a reasonable number of data for supervised learning and shows remarkable generalization power. We also introduce methods for intuitive and consistent image editing and restyling, e.g., by positioning, rotating or resizing individual objects in a scene. Our method integrates well within various image creation workflows and enables a richer set of applications compared to previous approaches.
1 Introduction
The integration of user-prompted conditioning signals from diverse modalities has been a major factor in the recent success of image generation models [23, 53, 39, 55], enabling precise control, expanding creative possibilities, and narrowing the gap between human intent and generated content. Yet, there is still room to make controllable image synthesis of multi-object scenes more intuitive and better integrated with conventional design and composition workflows. In a standard pipeline for photorealistic rendering, designers generally follow a bottom-up approach: first defining the three-dimensional structure of the environment, then introducing objects and their appearance before rendering from different viewpoints and, optionally, refining details through local edits or relighting. However, existing guiding mechanisms for generative image models predominantly rely on text [53, 39, 55] or image [53, 38, 78, 75] representations, which do not inherently preserve consistent, unambiguous, and well-defined 3D structures.
The dominant approach of generative image modeling through text guidance, while convenient, also makes it difficult to convey complex compositional structures [43], as it can be cumbersome to capture nuanced spatial and geometric relations through text alone. Such relations are especially prominent in multi-object environments such as indoor scenes. Consequently, approaches leveraging structural controls, such as bounding box layouts, keypoints or semantic maps [26, 53, 47], have emerged to provide more explicit control over composition. Key to their adoption has been their efficient integration into pretrained diffusion models via the training of lightweight adapter modules [78, 38, 75], that preserve prior generative capabilities, while enabling richer and more intuitive controls. Attempts aimed at providing 3D-aware control for image synthesis and editing have gained popularity [7, 73, 41, 58, 9, 71], but typically rely on panorama images, video datasets or conditioning through depth maps that are hard to acquire and offer limited flexibility.
Overall, and as illustrated in Figure 1, the nature of these inputs is inherently associated with critical limitations, as they discard crucial spatial information such as viewpoint, object orientations, or occlusions in nested arrangements. As a result, existing representations struggle to accurately account for cases where objects are placed within others—such as books stored inside a shelf—despite such configurations being highly prevalent in real-world scenes. They also encode geometric information in a highly viewpoint-dependent manner, causing inconsistencies in how spatial structures are perceived across views. Finally, conditioning should ideally encode 3D context in a comprehensive manner. This means, for instance, capturing the influence of off-screen elements, such as lighting from a window outside the frame, but also maintaining stylistic and functional coherence across the entire scene.
In this context, we propose to inject simple, explicit 3D geometric information as guidance for single-view generative models. Associated challenges are numerous in light of the limited availability of image data with accurate 3D semantic annotations [52, 11, 76], and the lack of established solution to encode 3D inputs as additional control to a pretrained diffusion backbone. From a technical perspective, incorporating conditioning in a form that is different from text or image inputs is indeed highly non-trivial. As a result, our work aims to introduce the representational and architectural components as well as the training dynamics to augment pretrained text-to-image models with guidance capabilities from parametric 3D semantic layouts. More precisely, our main contributions are threefold:
  1. 1.
    A parametric conditioning representation, based on semantic bounding boxes, which, for the first time, allows for 3D-informed image synthesis and editing, maintaining consistent structures across views and without relying on depth estimators, multi-view or panorama images.
  2. 2.
    An adapter architecture that establishes a tight relation between inputs in the 3D domain and the image target, while being compatible with pre-existing conditioning modalities.
  3. 3.
    An efficient training framework that enables camera control as well as geometric and free-form semantic object-level guidance in 3D.
We demonstrate the effectiveness of our method in a wide range of experimental settings, highlighting multiple advantages over current approaches. We additionally introduce a novel evaluation methodology that allows to evaluate the adherence to object-level conditioning, and use it to assess different image synthesis methods.
2 Related Work
”A square bedroom with a king-size bed, flanked by two nightstands. A wardrobe, approximately the same width as the bed, stands in front of it, with decorative items inside. Next to one of the nightstands, there is a large Alocasia plant. A window is positioned in front of the plant and beside the wardrobe. The plant is at the forefront, partially obstructing the view.”
(a) Text Description
(b) Typed 2D Bounding Boxes
(c) Semantic & Depth Maps
(d) Ours
Figure 1: Comparison of high-level conditioning input representations for describing a 3D scene. Relying solely on text descriptions (a) can make it difficult to convey complex spatial relations. Conditioning via 2D bounding boxes (b) can lead to ambiguity in perspective and does not account for out-of-bound objects. Semantic and depth maps rendered from 3D bounding boxes (c), as introduced in recent work [73], cannot always handle occluded and nested items while being limited to objects typed from a fixed set of categories. In contrast, our layout representation encodes a comprehensive 3D scene structure that is consistent across camera views and supports object-level semantic captioning and direct manipulation of position and orientation in 3D space.
3D-Aware Content Creation
Recent advances in deep generative modeling [18, 62, 22, 46, 28] have facilitated the emergence of powerful methods for user-driven content synthesis. In particular, score-based models have been employed to create realistic images [53, 55, 39], videos [61, 19, 48] or 3D assets [31, 30] of unprecedented quality. Despite these capabilities, achieving 3D-aware image synthesis and editing, i.e., that naturally incorporate the three-dimensional structure of the underlying scene, remains challenging. This task is typically tackled by leveraging multi-view [69, 4, 13, 12, 64] or multi-frame [71, 37] image datasets, or by using depth maps, which are hard to acquire and manipulate in real-life settings, as additional conditioning inputs [72, 41, 67]. 3D-aware approaches have notably demonstrated advantages over those relying on traditional text-to-image models for 3D scene generation [20, 9, 58]. However, these typically rely on costly score distillation [49] from the 2D prior [59, 6, 73], limiting their applicability.
Layout-Guided Image Synthesis
A prolific line of work has been incorporating structured spatial information to control image generation. Notably, compositional synthesis can be achieved from the guidance of 2D semantic bounding boxes [26, 53, 74, 79, 5] or segmentation maps [53, 68]. Closely related to our work, SceneCraft2D [73] renders 3D indoor layouts from any camera viewpoint by projecting object bounding boxes to depth and semantic maps, that are used as conditions to train respective ControlNet [78] modules. Other methods like ControlRoom3D [58] or Ctrl-Room [9] further rely on panorama images to render 360° consistent views of rooms. Recently, Build-A-Scene [7] proposes a training-free approach based on attention guidance [5] to perform 3D layout control from a depth-conditioned prior [2]. While convenient, this method does not scale to complex 3D layouts featuring more than a few objects. One common limitation of these approaches is that their conditioning input ultimately lies in the 2D space, making it difficult to disentangle objects in complex arrangements or to account for unseen contextual items from the 3D environment.
Adapters for Diffusion Priors
Adapters are lightweight learnable modules to be plugged into pretrained diffusion models. Notably, low-rank adapters [21] have been widely used [35] to efficiently perform customized domain adaptation, while structure adapters [78, 38] have enabled to control image generation with various 2D conditions that are spatially aligned with the content of the target image. More related to our work in its methodology, IP-Adapter [75] augments text-to-image models with image guidance, allowing to finely control the style and appearance of the synthesized content. This is done by learning additional cross-attention [65] weights, linking image embeddings extracted by pretrained foundation models [51, 40] with visual features from the diffusion backbone. In contrast, we propose to learn our 3D semantic layout encoder and attention-based adaptation module jointly.
Encoding Semantic 3D Layouts
Previous lines of work have proposed a range of representations based on graphs [77, 42, 25] to embed 3D arrangements in a way that can capture the relations between interacting elements. Concurrently, methods defining layouts as unordered sequences of objects, to be encoded by attention-based architectures, have shown compelling results in the context of 3D scene synthesis [44, 70, 63, 36].
Our method broadens the scope of existing control mechanisms and is the first to enable comprehensive control of both object semantics but also the underlying 3D geometric structure in a single unified framework.
3 Method
In this section, we introduce our framework, which is summarized in Figure 2. Our ultimate goal is to provide 3D layout control to pretrained text-to-image (T2I) models. Unlike previous work [73], we use an explicit, view-independent 3D representation and jointly learn feature extractors and adapter modules in a single training experiment.
Figure 2: LACONIC Architecture & Pipeline Overview. From an input 3D layout $\mathcal{S}$ and camera pose $\mathcal{C}$, trainable modules embeds geometric and semantic properties of individual objects to guide a text-to-image diffusion prior in denoising a target rendering. Camera control is enabled by expressing spatial features $(p,R)$ from the input object 3D bounding boxes $\mathcal{O}$ in the coordinate system defined by the target viewpoint. Resulting objects and, optionally, a room’s floor plan $\mathcal{F}$, are embedded by dedicated modules, yielding a set of representations $\mathcal{T}$ processed by a transformer-based encoder. Output sequence of embeddings is subsequently integrated as additional conditioning input to the pretrained T2I backbone via decoupled cross-attention [75].
3.1 Preliminaries
Our framework leverages text-to-image diffusion models (DM). We introduce the key associated generative and conditioning mechanisms that are relevant for our approach, and which we build upon below.
T2I Diffusion Models
T2I diffusion models are trained to denoise data samples $x$, perturbed by Gaussian noise $\epsilon$ across multiple timesteps $t$, given their associated text caption $c$ and, optionally, a task-specific condition $y$. This is done by parameterizing a neural network $\epsilon_{\theta}$ to predict the noise residual in $x_{t}$ following the simple objective:
$\mathcal{L}_{\text{DM}}=\mathbb{E}_{x,c,y,\epsilon\sim\mathcal{N}(0,I),t}\Bigl{[}\lVert\epsilon-\epsilon_{\theta}(x_{t},t,c,y)\rVert_{2}^{2}\Bigr{]}.$ (1)
Central to our approach, this learning objective can be used to incorporate diverse conditioning controls $y$ through fine-tuning experiments, leveraging a pretrained, frozen T2I backbone alongside additional trainable components.
Cross-Attention Conditioning
Input signals to control text-to-image generative models are learned through the use of cross-attention layers within the architecture. We introduce here the established mechanisms and notations that intertwine with our own conditioning approach, described in Section 3.3. In particular, given a tokenized text caption $c=[c_{1},c_{2},\ldots,c_{M}]$, a pretrained foundation model like CLIP [51] computes a sequence representation $\tau_{\theta}(c)\in\mathbb{R}^{M\times d_{\tau}}$ where $d_{\tau}$ is the text embedding size. Resulting key $K^{c}$ and value $V^{c}$, of hidden dimension $d$, are obtained from respective learnable linear projections. Similarly, the intermediate $h\times w$ feature map with $d_{\text{img}}$ channels from the image backbone, that is typically implemented as a UNet [54] with residual blocks, is flattened and projected to the query $Q$. New hidden state $H^{c}\in\mathbb{R}^{hw\times d}$ incorporating the text condition $c$ is finally computed via dot-product attention [65]. In practice, this mechanism is applied in a multi-headed fashion and at various feature resolutions within the conditional image denoiser $\epsilon_{\theta}$.
3.2 Semantic Layout Representation
At the heart of our method lies the additional conditioning input $y$ that we define as an intuitive and explicit proxy for furnished 3D layouts. As described in this section, individual objects are encoded from their high-level semantic and spatial properties using a parametric bounding box representation in the 3D space, as described below.
3D Scene Parameterization
We represent a 3D scene $\mathcal{S}$ as an unordered set of $N$ objects $\mathcal{O}=\{o_{1},\ldots,o_{N}\}$ and, optionally, an indoor floor plan modeled by a point cloud $\mathcal{F}\in\mathbb{R}^{P\times 3}$, where $P$ points are sampled along its boundary contours. We make $\mathcal{F}$ optional because (i) 3D layout datasets often lack detailed structural data, like floors, ceilings and walls, and (ii) we empirically found that 3D objects alone can implicitly reflect the scene’s structure via their spatial arrangements, given that they are commonly positioned on floors or aligned against walls. We represent each object by its “semantic 3D bounding box”: $o_{i}=(p_{i},d_{i},R_{i},s_{i})$. Here $p_{i}\in\mathbb{R}^{3}$ is the object’s center position in the world coordinate system, $d_{i}\in\mathbb{R}^{3}$ is the size along each dimension, and $R_{i}\in\mathbb{R}^{3\times 3}$ is the rotation matrix. Similar to the global scene caption $c$, the object-level semantic description $s_{i}=[s_{i}^{1},s_{i}^{2},\ldots,s_{i}^{M}]$ is processed to $M$ tokens from natural language. Those design choices are motivated by the recent success of similar lightweight representations, which have demonstrated their expressiveness in related 3D scene generative tasks [44, 36, 63, 70].
Camera Viewpoint
The camera $\mathcal{C}$ from which to render the 3D layout is represented by its extrinsic parameters, i.e., position $p_{\mathcal{C}}\in\mathbb{R}^{3}$ and rotation $R_{\mathcal{C}}\in\mathbb{R}^{3\times 3}$. Note that, for the purpose of our general methodology, we assume that the image samples $x$ from the data distribution are rendered using consistent camera intrinsics.
3.3 Adapter Architecture
In this section, we introduce the trainable modules and mechanisms designed to (i) capture meaningful representations from the scene conditioning input and (ii) incorporate them to the T2I backbone, which, as described in Section 3.1, is conditioned through cross-attention modules and remains frozen to fully retain its original capabilities.
3D Layout Encoder
The part of our architecture that is responsible for encoding the input 3D layout representation is inspired by previous work [70, 36, 44]. Notably, individual objects $o_{i}$ are handled by a common module that embeds scalar spatial features through sinusoidal positional encoding [65] and dense layers, while encoding the semantic caption using the pretrained text encoder backbone $\tau_{\theta}$. Concatenation of the resulting attributes produces per-object tokens $\mathcal{T}_{o_{i}}$. The optional indoor floor plan is encoded by a PointNet [50] module, leading to an individual token $\mathcal{T}_{\mathcal{F}}$. The sequence defined by the embedded tokens is passed to a transformer encoder computing new representations. Subsequently and following previous work [75], a dense unit shared between tokens outputs representations of dimension $d_{\tau}$, which is consistent with the global text embeddings, and is followed by Layer Normalization [1].
Decoupled Cross-Attention
Figure 3: Iterative Scene Editing Results. From left to right: given an input semantic 3D layout and camera viewpoint (a), we render the scene both without (b) and with (c) global text prompt conditioning. Then, individual objects are subsequently moved (d), resized (e) and re-captioned (f). Results highlight the ability of our approach to additively perform local manipulations and generalize to out-of-distribution concepts through the use of the global text conditioning, while maintaining a consistent 3D structure. Remarkably, we observe that semantic concepts from the global text prompt are represented on relevant objects in the generated image, e.g., patterns on wallpapers or bedsheets, art in wall frames, and do not leak to e.g., ceilings or floors. Interestingly, removing a window has an impact on the global illumination of the scene, which would have not been possible by editing local objects via e.g., 2D image inpainting [34].
The unordered sequence of scene representations $\hat{\mathcal{T}}=\{\hat{\mathcal{T}}_{\mathcal{F}},\hat{\mathcal{T}}_{o_{i}},\ldots,\hat{\mathcal{T}}_{o_{N}}\}$ output by the 3D layout encoder is intuitively ideal to establish the individual visual contributions of each object within the image backbone through cross-attention conditioning. It is projected accordingly to key $K^{y}$ and value $V^{y}$ both in $\mathbb{R}^{(N+1)\times d}$ by additional trainable linear projections. The hidden state associated to the input 3D layout is computed similarly to the one $H^{c}$ from the text condition, via dot-product attention with the same query $Q$ obtained from image feature maps:
$H^{y}=\text{softmax}\Biggl{(}\frac{Q{(K^{y})}^{\top}}{\sqrt{d}}\Biggr{)}\cdot V^{y},\quad H^{y}\in\mathbb{R}^{hw\times d}$ (2)
Following the decoupled cross-attention methodology [75], the final hidden state incorporating both the global text $c$ and 3D layout $y$ conditions is obtained via a weighted sum:
$H=H^{c}+\gamma H^{y},\quad H\in\mathbb{R}^{hw\times d}$ (3)
where $\gamma$ is a scalar controlling the strength of the scene control with respect to the global caption one. It is finally projected back to be added to image feature maps of the diffusion backbone’s residual blocks.
3.4 Supervised Training
At training time, we assume to be given a collection of scenes, from which to extract structural and semantic 3D bounding box representations, as described in Section 3.2. Each scene $\mathcal{S}$ is associated to a rendering image $x_{0}$ and corresponding camera pose $\mathcal{C}$. In this section, we describe the training dynamics adopted to efficiently map our additional guidance signal to the target scene image, enabling free camera control and object-level, open-vocabulary semantic conditioning.
Camera Viewpoint Transformation
One important challenge associated to our method is that our conditioning input is defined in a three-dimensional world coordinate system, while the generation target lies in the image domain. As a result, we introduce an explicit mechanism to handle the crucial translation between the 3D input representation and underlying output image. Importantly, the information of the input viewpoint from which $\mathcal{S}$ is rendered is directly integrated into the object spatial representations, by expressing them in the 3D coordinate system defined by the camera $\mathcal{C}$. More precisely, we apply the following series of transformations:
$p_{i}^{\mathcal{C}}=R_{\mathcal{C}}^{\top}(p_{i}-p_{\mathcal{C}}),\quad R_{i}^{\mathcal{C}}=R_{\mathcal{C}}^{\top}R_{i},\quad\forall o_{i}\in\mathcal{O}$ (4)
We found this reframing mechanism of the 3D input signal to the 2D image domain to be key in order to efficiently leverage the pretrained diffusion model. It can be performed in closed form on the fly, without requiring the network to learn such complex mapping between domains. When it is provided, a similar transformation is performed with the floor point cloud input $\mathcal{F}$. In practice, object rotation features $R_{i}^{\mathcal{C}}$ are further mapped to a continuous representation, following [81]. The resulting scene serves as the additional conditioning input $y=\mathcal{S}^{\mathcal{C}}$ to guide the pretrained T2I model $\epsilon_{\theta}$ in denoising the target image $x_{0}$ consistent with the 3D layout structure and camera view, following the training objective from Equation (1).
Conditioning Dynamics
We adopt classifier-free guidance [17] training by randomly dropping the 3D layout input $y$, which, at each iteration, is set to $\emptyset$ with probability $\mathbf{p}_{\text{drop}}$, so that the denoiser network models both the layout-conditional and unconditional image densities. Additionally, image targets $x$ do not need to be associated to a global textual description and, as a result, $c$ is always obtained from an empty prompt during training. Unlike previous work [73], which relies on one-hot category representations, our method conditions individual objects with free-form text captions $s_{i}$. If not directly part of the dataset annotations, caption supervision can be derived from a vision-language model [29, 24] applied to an image showcasing the object of interest.
3.5 Application Scenarios
As illustrated in Figure LABEL:fig:teaser, we provide a broad overview of the range of generative and editing capabilities enabled by our trained model at test time. Here, we consider an initial input scene $\mathcal{S}_{0}$, which can be conveniently user-provided by disposing, sizing and describing a collection of objects.
Generating Structurally-Consistent Views
Although our method is designed to be trained on single-view datasets, the comprehensive 3D context provided by our input representation allows synthesizing multiple views of a given scene, sharing a consistent 3D structure. More formally, from a set of target camera viewpoints $\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\}$, the corresponding inputs $\mathcal{S}_{0}^{\mathcal{C}_{i}}$ conditioning the trained model can be obtained by applying the space transformation logic from Equation (4).
Text-Driven Scene Restyling
Remarkably, while our model has not been trained with global text caption $c$ supervision, its adapter design allows to fully leverage the prior knowledge from the T2I backbone, benefiting from strong out-of-distribution generalization to a wide range of semantic concepts. Notably, as described in Section 4.2, strict adherence to the input scene and viewpoint can be relaxed by lowering $\gamma$ in Equation (3), thus interpolating structural variations matching the global input description.
Object Attribute-Level Scene Editing
The initial input scene $\mathcal{S}_{0}$ can be iteratively edited based on user preferences, by modifying parts of the objects in $\mathcal{O}_{0}$ or adjusting its items set through additions or removals. Importantly, individual objects can be manipulated with per-attribute granularity i.e., by individually and independently adjusting their size, positioning or semantic features. This shows practical advantages over previous approaches [41] whose attempts to solve 3D-aware image editing required to edit conditioning depth maps in the pixel space.
4 Experiments
Figure 4: 3D layout-guided image synthesis baseline comparisons. Our method produces more detailed and natural images compared to baseline approaches. Methods leveraging our 3D layout encoder (DM-FS & ours) better represent the guiding layout, while our adapter-based approach additionally shows advantages at producing higher quality images. We also observe in text-driven scenarios (e–h) that compared to SceneCraft [73], LACONIC better blends conditioning inputs coming from different sources, adhering both to elements from the 3D layout and the semantic text caption conditionings.
In this section, we showcase our method’s capabilities at generating realistic images that incorporate the desired semantic and geometric input information. To this end, we conduct various experiments comparing it to baseline approaches on a range of standard metrics. We also introduce a novel metric that captures the alignment between the 3D layout prompt and the rendered image on the object level.
4.1 3D Layout-Guided Image Synthesis
Datasets
We follow previous work [73] and use HyperSim [52] indoor scene dataset to extract semantic 3D bounding box layouts as well as camera poses with their associated photorealistic renderings. As HyperSim only features $326$ unique 3D layouts, paired to a total of $24{,}383$ images, methods taking as input the comprehensive 3D layout associated to each image sample, such as ours, may tend to memorize individual fixed scenes from the dataset, making it difficult to perform the complex out-of-distribution manipulations of individual objects that our framework enables. In response, and in order to provide additional qualitative results showcasing the full range of our model’s capabilities, we also gather a custom dataset featuring $72{,}000$ bedroom 3D scene layouts, each being paired to a single, high-quality image and viewpoint. See supplementary for additional details and statistics on the used datasets.
Baselines
We ensure a relevant and fair comparison by competing against SceneCraft [73], a recent baseline which, in the like of our method, (i) proposes to synthesize 2D images from 3D layout controls, (ii) is based on supervised training on single-view, non-panoramic images and (iii) is an adapter-based approach, enabling text-driven synthesis at test time. Additionally, we evaluate the impact of using our adapter approach instead of training a layout-conditioned Diffusion Model From Scratch (DM-FS) on layout-image pairs by introducing a dedicated baseline that retains the 3D layout conditioning encoder from our methodology.
Implementation
Figure 5: Impact of the adapter strength on generated images. From a text caption $c=$ ”A Retro Synthwave Indoor Bedroom”, adherence with the input 3D layout improves accordingly with the parameter $\gamma$ from Equation (3), achieving strict control at higher scales. We can observe that even at lower scales (c) the control enforces the geometric and semantic plausibility of the synthesized image.
Throughout our experiments, we employ Stable Diffusion v1.5 [53] as the pretrained text-to-image diffusion backbone upon which our adapter network is trained. We train the DM-FS baseline using a similar but downsized UNet architecture, to account for the scale of the dataset. SceneCraft [73] is used with its pretrained weights and default parameters from the official implementation. All the compared methods in quantitative evaluations and side-by-side qualitative comparisons are trained on the same HyperSim [52] subset. We provide comprehensive implementation details as well as training and inference settings in the supplementary materials.
Metrics
[…]
Scene Object CLIP score (SOC)
We also evaluate different methods on a new metric that we call Scene Object CLIP score (SOC). Our score is motivated by the fact that most established metrics for evaluating image generative models primarily focus on image quality, diversity, and prompt adherence, rather than spatial or semantic alignment at the object level. Instead, SOC explicitly aims at measuring the alignment of individual objects in the conditioning layout with the corresponding locations in the synthesized image.
Figure 6: Illustration of our metric’s object selection mechanism. Obstructed and out-of-bounds objects are identified from their 2D and 3D bounding box annotations and filtered out from the evaluation set.
Our metric takes as input a synthesized image, a guiding 3D semantic layout and 2D object bounding box $\mathcal{B}_{\text{2D}}$ annotations from the ground truth image. To compute SOC, we proceed as follows: for each object in the 3D layout, we (i) project its 3D bounding box on the rendered image using the camera $\mathcal{C}$ extrinsic and intrinsic parameters, (ii) derive the corresponding 2D enclosing bounding box $\mathcal{B}_{\text{3D}}$, where $\mathcal{B}_{\text{2D}}\subset\mathcal{B}_{\text{3D}}$ and (iii) define a minimum threshold on the intersection area ratio $\mathcal{B}_{\text{2D}}/\mathcal{B}_{\text{3D}}$ for an object to be considered sufficiently visible. Finally, we crop the selected items from their enclosing bounding box $\mathcal{B}_{\text{3D}}$, obtaining a collection of per-object images $x_{o_{i}}$. These steps are meant to only keep objects in 3D space which are expected to be sufficiently visible in the given 2D image, as illustrated in Figure 6.
Since each visible object $o_{i}$ is associated with a text description, we can compute a semantic correlation between the object’s text caption $s_{i}$ and its crop image $x_{o_{i}}$. For this last step, we follow [15] among others, and use the pretrained CLIP [51] model. We provide details on our Scene Object Clip score (SOC) in the supplementary materials.
Table 1: Quantitative experiment results on 3D layout-guided images synthesis. We compare our method against other learning-based approaches that we outperform on most evaluation metrics, both with and without providing a global text caption $c$.
Methods FID $\downarrow$ KID $\downarrow$ IS $\uparrow$ CS $\uparrow$ SOC $\uparrow$
SceneCraft [73] w/o text prompt 39.36 28.26 7.72 17.59
DM-FS 15.83 7.29 8.69 18.22
ours 9.50 3.44 9.74 18.36
SceneCraft [73] w/ text prompt 27.69 15.21 14.55 19.75 17.40
ours 10.12 3.91 10.60 19.74 18.39
Results
Our main quantitative results are summarized in Table 1. We report results both with and without text prompting. In the latter case our approach achieves state-of-the-art results across all metrics and outperforms the very recent baseline SceneCraft [73] by a significant margin. When using text conditioning (the captions provided in HyperSim [52]), our method still outperforms SceneCraft [73] on most metrics and in several cases by a significant margin. We note that our approach is able to capture the HyperSim data distribution significantly better compared to SceneCraft [73]. However, due to the relative limited scope of 3D structures in that dataset, our approach demonstrates somewhat less variety in text-driven scenarios as indicated by a lower Inception Score. Importantly, our adapter-based approach outperforms the DM-FS approach on all metrics, while enabling text-conditioned synthesis, thus validating the importance of using a rich pretrained T2I model. We provide qualitative comparisons on the 3D layout-guided image synthesis task in Figure 4. Notably, we observe that our approach produces more natural images, featuring fewer visual artifacts compared to baseline methods, while being more faithful to the input 3D layout and viewpoint.
4.2 Additional Qualitative Results
We provide additional generation results, showcasing the intuitive and iterative editing capabilities of our framework as well as its ability to leverage the pretrained text-to-image backbone for out-of-distribution generalization in Figure LABEL:fig:teaser and Figure 3. Overall, these results highlight that our model can capture nuanced individual object relations within structured scenes. Moreover, our approach enables manipulations in 3D space, and provides wholistic scene modeling, accounting, e.g., for off-screen objects. Finally, results reported in Figure 5 show how introducing our adapter control, even at low scale $\gamma$, allows synthesizing 3D-aware images that are more geometrically and semantically sound than base text-to-image outputs, while maintaining adherence with the textual description. Additional qualitative results highlighting the power of our approach can be found in the supplementary materials.
5 Conclusion, Limitations & Future Work
In this paper, we introduced LACONIC, a novel 3D layout-conditional image generative framework that is distinct in both representation and architecture, expanding the creative capabilities in 3D-aware image synthesis and editing. By efficiently mapping the underlying 3D representations to the image domain and integrating them to pretrained text-to-image model, our method introduces fine-grained structural and semantic control, enhancing both user interactivity and the expressiveness and coherence of generated content.
While powerful and capable of generalization to unseen concepts, our model’s output domain is still bounded by the training data distribution. As a result, a model trained on bedroom layout data will be unlikely to synthesize e.g., plausible kitchens while allowing fine-grained control. We believe that our model would be ideal to derive continuous 3D scene representations, such as NeRFs, that finely adhere to user-defined structures from the conditioning input. We also believe that exploring ways to couple geometric and visual consistency of objects across changes in camera views is an important future research direction.
Supplementary Material
Figure 7: Additional image synthesis results (Right) from an input 3D layout and prompt (Left). Our method shows great adherence to both input 3D layout and caption prompt. Notably, LACONIC generates high quality results without needing prompt information (bottom center), with in-domain prompts (top right & bottom left) and with out-of-distribution prompts (top left/center & bottom right).
Appendix A Implementation
In this section, we provide comprehensive implementation details on our 3D layout adapter architecture, pretrained text-to-image diffusion backbone, as well as our training and inference settings. Unless explicitly specified, same settings are shared across the different datasets that we employed to conduct our experiments. We also detail the training and test configurations of the baseline methods.
A.1 Network Architecture
A.1.1 Pretrained Models
We employ the established and widely adopted Stable Diffusion [53] v1.5 to implement the conditional UNet image denoiser $\epsilon_{\theta}$ and CLIP [51] text prompt encoder $\tau_{\theta}$, using the pretrained weights and implementation [^1] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 from the HuggingFace Diffusers [66] library. In our framework, the text encoder is used to encode both the global image caption $c$ and the object-level semantic descriptions $s$.
A.1.2 3D Layout Encoder
We detail the individual trainable modules that, building upon previous work [70, 36, 44], are used to embed the input semantic 3D layout $\mathcal{S}$, as illustrated in Figure 2 and described in Section 3.3 in the main paper.
Shared Object Encoder
Parametric object bounding box representation $o_{i}$ are embedded by a common module. More precisely, each scalar defining individual object spatial attributes $(p_{i},R_{i},d_{i})$ is projected to a higher dimensional vector using fixed, sinusoidal positional encoding [65] with $32$ frequencies. We follow [70] and apply:
$PE(k)=\{\sin(128^{j/31}k),\cos(128^{j/31}k)\}_{j=0}^{31}\in\mathbb{R}^{64}$ (5)
Consequently, object’s position $p_{i}$ and dimension $d_{i}$ are encoded to $192$-dimensional attributes. Importantly, the rotation matrix $R_{i}$ is first expressed in a continuous representation, following the recommendations from [81], leading to a vector in $\mathbb{R}^{6}$ from which individual scalars are also processed by Equation (5). This produces a representation in $\mathbb{R}^{384}$ that is subsequently mapped by a linear layer in $\mathbb{R}^{192}$. The semantic embedding $\tau(s_{i})$ is the end-of-sequence output token from the pretrained text encoder in $\mathbb{R}^{d_{\tau}}$, with $d_{\tau}=768$, and is further projected to a $192$-dimensional representation by an MLP featuring a single hidden unit of dimension $384$ with LeakyReLU activations. The mentioned encoded object attributes are finally concatenated, yielding an individual token $\mathcal{T}_{o_{i}}$ of dimension $4\times 192=d_{\tau}$ for each of the $N$ objects in the scene.
Floor Plan Encoder
Following prior work [70, 36], we encode the optional scene’s floor plan $\mathcal{F}$ by leveraging a popular implementation [^2] https://github.com/fxia22/pointnet.pytorch of the PointNet [50] model, applied on a point cloud representation obtained by sampling $P=100$ three-dimensional points, that are evenly spaced along the room’s boundaries. A single floor token $\mathcal{T}_{\mathcal{F}_{i}}\in\mathbb{R}^{d_{\tau}}$ is subsequently obtained by projecting the $1024$-dimensional module’s output.
Transformer Encoder
New token representations $\hat{\mathcal{T}}$ are established by a Transformer encoder that follows the seminal paper [65] by using the implementation from the PyTorch [45] library. In practice, we use $4$ encoder layers with $6$ attention heads and hidden dimensions of size $512$. Since the scene is defined as an unordered sequence of its objects and floor representations, we do not additionally encode the position of individual tokens in $\mathcal{T}$, and perform zero-padding to handle scenes with different number of objects $N$. In practice, we consider a maximum number of $50$ objects. Following [75], the output token embeddings are each passed to a shared MLP that preserves the token dimension $d_{\tau}$, implemented with a hidden unit of size $768$, GELU activation [14], and followed by Layer Normalization [1].
A.1.3 Cross-Attention Layers
As introduced in Section 3.3 in the main paper, the scene conditioning sequence $\hat{\mathcal{T}}$ outputted by the 3D layout encoder is mapped to associated key $K^{y}$ and value $V^{y}$ by introducing respective learnable dense projection matrices $W_{K}\in\mathbb{R}^{d\times d_{\tau}}$ and $W_{V}\in\mathbb{R}^{d\times d_{\tau}}$. In practice, we follow the implementation [^3] https://github.com/tencent-ailab/IP-Adapter from IP-Adapter [75] to introduce the decoupled cross-attention mechanism within residual blocks from the pretrained Stable Diffusion UNet backbone.
A.2 Datasets
In this section, we introduce the datasets that we used to conduct our experimental evaluations and introduce any relevant dataset-dependent mechanisms implemented to adapt our work to their specific features or available annotations.
A.2.1 HyperSim Dataset
Features
Following previous work [73], we leverage indoor scenes and photorealistic renderings from HyperSim [52] to construct a collection of semantic 3D bounding box layouts with camera poses $y$ and image $x$ pairs. The dataset originally features $461$ scenes from diverse types of indoor environments, and $77{,}400$ rendered images. We follow SceneCraft [73] and leverage a filtered version [^4] https://huggingface.co/datasets/gzzyyxy/layout_diffusion_hypersim of the dataset proposed in [58], discarding unbounded and scenes with excessively large scales. This results in $323$ unique scenes, that are associated to $24{,}383$ images annotated by a global text caption extracted by a pretrained foundation model. Each layout features 3D bounding box annotations that are typed according to the standard NYU40 [60] semantic classes. Each rendering is also associated to an object instance map, that we leverage in the context of our SOC metric implementation, as detailed in Section A.5. We conducted foundational statistical analysis of the 3D layouts, establishing that scenes contain an average of $121$ objects, with a median of $54$ and that the dataset demonstrates a large diversity in terms of structural complexity, which is challenging in the context of our work and in light of the limited number of available samples.
Experimental Setup
We describe here the dataset-specific mechanisms that we implemented to run experiments on HyperSim layouts. Importantly, a proportion of scenes from the dataset feature hundreds of objects, exceeding the maximum length fixed to $50$ for our sequence-based representation used for the model conditioning. As a result, we leverage the instance semantic map of the target rendering to prioritize including visible objects in the conditioning sequence. In case the number of visible objects still exceeds the maximum length, we identify the most meaningful objects based on their bounding box dimensions. As the HyperSim dataset does not provide the floor information in a straightforward manner, we set $\mathcal{F}=\emptyset$ in our experiments. Object-level semantic information are also limited to their categorical label among NYU40 classes, from which we apply $\tau_{\theta}(s_{i})$ as described in our general methodology.
A.2.2 Custom Bedroom Dataset
As described in the main paper in Section 4.1, the limited number of available HyperSim layouts, that are passed as conditioning input, makes it easier for our model to learn a mapping between the global scenes and specific camera poses and target renderings, while overlooking the contributions of individual objects. In response and for the purpose of our qualitative evaluations, we gathered a custom dataset of $76{,}000$ human-designed 3D indoor bedroom layouts, in which each layout is associated to a single rendering. The dataset also includes floor information, represented as a 3D point cloud. Finally, ground-truth object-level semantic descriptions $s_{i}$ are extracted using a LLaVA model [29] from their 2D rendering using the following instruction: ”Describe this object concisely. Do not analyse the background, only the object”. This level of annotations allows to implement all the building blocks from our general methodology.
A.3 Training Protocol
To optimize our adapter network following a diffusion-based approach with frozen text-to-image denoiser, we follow standard DDPM [18] training with $\epsilon$-prediction objective, as formalized by Equation (1) in the main paper, and uniformly sampled noise timesteps $t\sim\mathcal{U}(0,1000)$. We randomly drop the conditioning input 3D layout $y$ with a rate $\mathbf{p}_{\text{drop}}=0.15$. We use the AdamW [32] optimizer, with weight decay coefficient $\lambda=0.01$ and learning rate $\eta=5\times 10^{-5}$, reached following an initial linear warmup phase during the first $200$ training iterations. Our checkpoints are trained for a total of $250$ epochs, with $\eta$ decreasing according to a cosine schedule. We leverage PyTorch Lightning [8] to distribute experiments across $3$ NVIDIA RTX 6000 GPUs, each handling a batch size of $10$ samples. Importantly, we experimented enabling automatic mixed precision during training, but empirically observed instability and inconsistent convergence.
A.4 Baselines
A.4.1 SceneCraft
SceneCraft [73] is a recent baseline method that proposes to tackle 3D layout-guided image synthesis, in which, similar to our approach, 3D layouts are defined as a collection of typed object 3D bounding boxes. To do so, it renders the input 3D layout from the target camera viewpoint to (i) a semantic map in which individual boxes are colored according to their one-hot semantic category and (ii) depth maps that are automatically derived from the input geometry. From this pair of 2D input representations, SceneCraft trains dedicated ControlNet [78] modules in supervised experiments and in order to introduce the additional controls to a pretrained text-to-image backbone. In practice, we leverage the pretrained checkpoint [^5] https://huggingface.co/gzzyyxy/layout_diffusion_hypersim_prompt_one_hot_multi_control_bs32_epoch24 obtained by training the modules on the same HyperSim [52] subset that we used in our own experiments. It uses Stable Diffusion [53] v2.1 as the T2I prior. We employ official author implementation [^6] https://github.com/OrangeSodahub/SceneCraft with the default test-time hyperparameters.
A.4.2 Diffusion Model from Scratch
We describe below the main motivations and implementation details behind the Diffusion Model trained From Stracth (DM-FS) from our experimental evaluations.
Intuitions
An important design choice of our framework is to rely on a cross-attention-based adapter network, in order to augment existing T2I models with 3D layout guidance, instead of training dedicated generative models from scratch. The main intuitions behind this critical choice are that (i) considering the small scale of available image datasets featuring 3D layout annotations, achieving photorealistic generation results would be extremely challenging, especially on challenging camera views that are less represented in the training distribution and (ii) we could leverage prior knowledge from other domains to achieve powerful generalization, as shown throughout our experimental results in the main paper, which would not be possible otherwise. We believe that those insights validate the relevance of this baseline approach in the context of our contributions.
Implementation
To experimentally validate our intuitions, we propose to train jointly a randomly initialized diffusion UNet backbone and the 3D layout encoder detailed in Section A.1.2. The backbone is similar to the one of Stable Diffusion, implemented with Diffusers [66], and is guided through regular cross-attention conditioning [53] between the 3D scene sequence $\mathcal{S}$ and feature maps from the residual blocks. Considering the very little variety of image captions in the training datasets, it is not trained with additional text conditioning. The UNet is also downsized, compared to Stable Diffusion, to account for the smaller scale of the training set. Importantly, we stack $4$ downsampling and upsampling residual blocks with $128$, $256$, $256$ and $384$ output channels. The $2$ bottleneck blocks, that operate at smaller resolutions, are augmented with cross-attention with $8$ heads. For the 3D layout encoder, we keep the same parameters as those described in Section A.1.2.
A.5 Evaluation Metric
In Section 4.1 of our main paper, we introduce the Scene Object CLIP (SOC) score as a robust way to simultaneously assess that objects in the generated content (i) are correctly positioned and sized with respect to their spatial conditioning information and, at the same time, (ii) match their assigned semantic attributes.
Procedure
To do so, given a conditioning semantic 3D layout $\mathcal{S}$ and camera pose $\mathcal{C}$, its associated ground truth image $x_{0}$ with 2D object $o_{i}$ bounding box annotations $\mathcal{B}_{\text{2D}}^{i}$ and, importantly, a synthesized image that we wish to evaluate, we identify the main visible objects composing the rendered scene. As 2D bounding boxes $\mathcal{B}_{\text{2D}}^{i}$ from the ground truth annotations alone do not assess that the corresponding objects $o_{i}$ are visible in their most part, we also leverage the 3D bounding box from the conditioning signal projected to image coordinate, from which we derive the enclosing 2D bounding box $\mathcal{B}_{\text{3D}}^{i}$. We notice that this second bounding fully contains the one from the image annotation, and compute the ratio $r_{a}$ of the area of $\mathcal{B}_{\text{2D}}^{i}$ over the one of the enclosing projected box $\mathcal{B}_{\text{3D}}^{i}$. Remarkably, this projection mechanism also handles objects that partially lie outside the image bounds. Based on this ratio, objects $o_{i}$ that are sufficiently visible are cropped at the location of $\mathcal{B}_{\text{3D}}^{i}$ in the synthesized image, yielding per-object images $x_{o_{i}}$. We finally compute a CLIP [51] correlation score between object images and their associated text semantic $s_{i}$.
Filtering Parameters
We rely on an object’s associated visible area ratio $r_{a}^{i}$ to determine if its crop image will be added to the evaluation base. In practice and to report the values in Table 1 in the main paper, we set a threshold value $\alpha=0.4$. Also note that tiny objects, that are widely represented in the HyperSim [52] dataset are initially filtered out. Indeed, objects whose instance semantic map spans less than $2$% of the image pixels are discarded. Finally, we also filter objects whose NYU40 semantic annotation is not meaningful, e.g., that are classified as ”other”. The metric is averaged over all the resulting object crops in the dataset.
Appendix B Additional Results
We provide additional qualitative results on text-driven, layout-guided image synthesis in Figure LABEL:fig:teaser. This highlights the ability of our method to adhere with the input semantic 3D layout while being capable of out-of-distribution generalizations to various unseen concepts. Additionally, we report in Figure 9 and Figure 10 supplemental comparisons between the DM-FS and SceneCraft [73] baselines, both with and without providing an input text as an additional condition. Our method consistently generates more realistic images, with better details as well as 3D layout and text prompt adherence. Motivating our 3D layout conditioning encoder, the DM-FS baseline appears to better respect the input 3D layout in comparison to SceneCraft [73]. We notice that both adapter-based approaches, ours and SceneCraft, leverage the pretrained T2I backbone, resulting in less artifacts in the generated images in comparison to DM-FS. Consequently, our method also show strong competitive advantages over SceneCraft and DM-FS when the number of objects in the scene is high thanks to the combination of the expressiveness of our 3D layout conditioning encoder coupled with the ability of the T2I backbone to generate detailed visual features.
Appendix C Limitations
Figure 8: Failure Cases and Limitations. In (a), some objects, such as the windows, are missing in the generated image and the floor shape, while complex, is not accurately respected. We can also observe visual stability issues, as shown in the result in (b), which features an inconsistent pattern on the wall resembling a misplaced frame above the bed. Finally, generated images can exhibit unnatural perspective, resulting in distorted floors and objects, as showcased in (c). This behavior may result from our assumption of consistent camera intrinsics across all data samples.
We provide failure cases and known issues in Figure 8. While showcasing state-of-the-art results on 3D layout adherence, our method can still lead to some inconsistencies in the generated results.
Appendix D Societal Impact
Controllable generative AI for 3D environments has significant implications for both individuals and industries. We believe our method to predominantly yield positive societal impact. By enabling fine-grained control over generation, our approach makes 3D-aware content creation more intuitive and accessible, empowering not only artists and designers but also individuals with no prior expertise in 3D modeling. It also benefits industries that rely on realistic environmental rendering, such as urban planning, architecture, and digital twins for cities. A key advantage of our method is its parameter efficiency, which reduces the computational cost of training—an important factor given the high environmental impact of generative models [33, 27]. However, these benefits come with challenges. Since our approach does not natively incorporate recent advances in generative model watermarking [10, 57], it could be misused for potentially deceptive applications. Additionally, by significantly lowering the barrier to content generation, it may contribute to rapid job displacement in traditional 3D design fields. Another concern is the potential for bias propagation, as image generative models, including semantic and style-driven ones like ours, may unintentionally reinforce stereotypes [80]—for example, by associating certain colors with gender, by e.g., synthesizing a bedroom as blue for ”boys” and pink for ”girls”. Ensuring the responsible development and use of such technologies, while addressing ethical concerns and mitigating biases, is crucial.
Figure 9: Additional 3D layout-guided image synthesis baseline comparisons (1/2). We can observe that our method produces more natural images and with respect to the input 3D layout.
Figure 10: Additional 3D layout-guided image synthesis baseline comparisons (2/2).